OpenStudy (anonymous):

Solve exactly. 8^(1-y) = 4^(y+2) y= ?

OpenStudy (anonymous):

Note that \(8 = 2^3\) and \(4=2^2\). So re-writing your equation using this form and taking the \(log_2\) of both sides will probably do nice things.

OpenStudy (anonymous):

Or just equating the exponents once you have both sides as a power of two amounts to the same thing.

OpenStudy (anonymous):

hmmm not following

OpenStudy (anonymous):

so 2^3(1-y) = 2^2(y+2)

OpenStudy (anonymous):

\[8 = 2^3\] \[\implies 8^{1-y} = (2^3)^{1-y} = 2^{3(1-y)}\] \[4 = 2^2\] \[\implies 4^{y+2} = (2^2)^{y+2} = 2^{2(y+2)}\] \[\therefore 8^{1-y} = 4^{y+2} \]\[\implies 2^{3(1-y)} = 2^{2(y+2)} \]\[\implies 3(1-y) = 2(y+2)\]

OpenStudy (anonymous):

That's right. You're on the track now.

OpenStudy (anonymous):

Take the natural log of both sides then take the exponents (1-y) and (y+2) down to the front of each term as factors then gather the y variables and factorise and simplify.

OpenStudy (anonymous):

and the 2's cancel out because you are using log 2?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Or just equating exponents.

OpenStudy (anonymous):

hmmm I get -1/5

OpenStudy (anonymous):

I get postive 1/5

OpenStudy (anonymous):

Double check your algebra.

OpenStudy (anonymous):

Oh wait. I should check mine. -1/5 is right.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

-0.2 or -1/5 is correct. :)

OpenStudy (anonymous):

lol was gonna say that what I keep getting